the following data are taken from alem corporation for the current year

material inventory beginning of year 22000 end of year 26000

WIP inventory beginning of year 36000 end of year 30000

F/G inventory beginning of year 18000 end of year 23000

putchase of R/M end of year 75000

direct labor 82000

indirect labor 15000

factory ins (OMOH) 9000

Dep on factory (OMOH) 11000

repair maintenance factory 4000

marketing expense 63000

general and admin expense 29000

income tax expense 30000

1.compute cost of goods mfg

2.prepare income statement assuming that 27000 units are sold at a price of $14 during current year


1
Expert's answer
2021-08-20T07:03:43-0400

a. Cost of goods manufactured

Cost of goods manufactured = Direct materials + direct labor+ manufacturing overhead + beginning WIP Inventory - Ending WIP inventory.

Direct materials{ (Opening Material Inventory + material purchased -End material inventory)

= 22000+ 75000- 26000 = 71000


Direct materials 71000

Direct Labor 82000

Manufacturing overheads:

Indirect Labor 15000

factory ins (OMOH) 9000

Dep on factory (OMOH) 11000

repair maintenance factory 4000

Total overhead costs 19000

Total manufacturing costs added =(71000+82000+19000)

=172000

Add: beginning WIP 36000

Less: Ending WIP (3000)

Cost of goods mfg =178000


b. Income statement

Revenue

Sales (27000 x 14) 378000

Cost of goods sold

cost of goods manufactured: 178000

Add: beginning finished goods: 18000

Cost of goods available: 196000

Less ending finished goods inv. (23000)

cost of goods sold 173000

Gross margin (sales-cost of goods sold) 205000

Less operating expense

marketing expense: 63000

General and Admin exp. 29000

Total operating exp. 92000

Income before tax (margin- oper. exp) 113000

Less income tax exp. (30000)

Net Income 83,000